Assume the following information:

Amount Per Unit
Sales $ 600,000 $ 40
Contribution margin $ 360,000 $ 24
Net operating income $ 240,000


If the selling price per unit increases by 6% and unit sales drop by 4%, then the best of estimate of the new net operating income is:
Answer #1

Current sales=($600,000/40)=15000 units

Hence new sales=(15000*(1-0.04)=14400 units

Contribution margin-Fixed cost=Net operating income

Fixed cost=(360,000-240,000)=$120,000

Sales (40*1.06)*14400 610560
Variable expenses(14400*24) 345600
Contribution margin 264960
Fixed expenses 120,000
Net operating income $144960

NOTE:Total fixed costs and variable cost per unit do not change with change in units